OpenStudy (atlchic):

Find the slope of the line through, (-2,-8) and (1,4).

OpenStudy (nottim):

goddarn, new quns keep popping up. i keep clicking wrong one

OpenStudy (nottim):

USe the slope formula: M=y/x=(Y2-Y1/X2-X1)

OpenStudy (nottim):

ok buddy?

OpenStudy (anonymous):

the slope is 4

OpenStudy (nottim):

YOU RUINED IT FAITH

OpenStudy (anonymous):

how??

OpenStudy (atlchic):

IM A GURL, SO 12/3

OpenStudy (nottim):

i wanted her to give me the answer, not you!

OpenStudy (nottim):

and what is 12 divided by 3?

OpenStudy (anonymous):

well i figured it out...

OpenStudy (nottim):

i wanted her to figure it out!

OpenStudy (anonymous):

WELLL I GOT IT FIRST

OpenStudy (nottim):

thats not the point of asking for help! (LOL)

OpenStudy (anonymous):

well i gave her help just very straight to the point help

OpenStudy (atlchic):

Write an equation of this line: Slope-5 & Y-Intercept--15

OpenStudy (nottim):

maybe ask that as a new question atlchic

OpenStudy (anonymous):

i just noticed that atlchic picture was flowers

OpenStudy (nottim):

can someone give me a shiny medal.

OpenStudy (atlchic):

go 2 my new question folks

OpenStudy (atlchic):

go 2 my new question folks